# Filling the gap in the IERS C01 polar motion series in 1858.9–1860.9

### v1

@@ Abstract
The IERS C01 Earth orientation parameters (EOP) series contains the longest reliable record of the Earth’s rotation. In particular, the polar motion (PM) series beginning from 1846 provides a basis for investigation of the long-term PM variations. However, the pole coordinate $Y_{p}$ in the IERS C01 PM series has a 2-year gap, which makes this series not completely evenly spaced. This paper presents the results of the first attempt to overcome this problem and discusses possible ways to fill this gap. Two novel approaches were considered for this purpose: deterministic astronomical model consisting of the bias and the Chandler and annual wobbles with linearly changing amplitudes, and statistical data-driven model based on the Singular Spectrum Analysis (SSA). Both methods were tested with various options to ensure robust and reliable results. The results obtained by the two methods generally agree within the $Y_{p}$ errors in the IERS C01 series, but the results obtained by the SSA approach can be considered preferable because it is based on a more complete PM model.
@@ 1 Introduction
Time series of the Earth orientation parameters (EOP) are the main source of information for investigation of the Earth’s rotation, in particular the motion of the Earth’s rotation axis in the terrestrial reference frame, which is also called polar motion (PM). The PM is a complex phenomenon, which consists of many variational components including trends and (quasi)periodical oscillations with periods from several hours to several decades. To study slow PM variations, long time series of the pole coordinates $X_{p}$ and $Y_{p}$ is needed. The longer the time interval covered by EOP series, the lower frequency variations can be detected and investigated.
The international standard of the EOP series is the combined series computed by the International Earth Rotation and Reference Systems Service (IERS) (Bizouard et al., 2019). The IERS C01 EOP series contains the longest PM series derived from astronomical observations lasting from 1846 to now111ftp://hpiers.obspm.fr/iers/eop/eopc01/eopc01.iau2000.1846-now. It is widely used for analysis of the decadal variations in Earth’s rotation (Rykhlova and Kurbasova, 1990; Kołaczek and Kosek, 1999; Gambis, 2000; Yatskiv, 2000; Höpfner, 2004; Guo et al., 2005; Malkin and Miller, 2010; Zotov, 2010; Miller, 2011; Seitz et al., 2012; Zotov and Bizouard, 2012; Beutler et al., 2020; Lopes et al., 2022; Zotov et al., 2022; Yamaguchi and Furuya, 2024).
@@
In this paper, for the first time, to our knowledge, we examine possible approaches to address the problem of filling the gap in the IERS C01 series. The challenge is to fill in the missing $Y_{p}$ values in the series while preserving its structure as much as possible. For this purpose, two methods of analysis were considered. The first is the use of a deterministic astronomical model, specially developed for this study and consisted of three components: bias, Chandler wobble (CW), and annual wobble (AW) (hereafter referred to as BCA model). The second data-driven statistical model has been constructed using the Singular Spectrum Analysis (SSA) (Golyandina et al., 2001).
The paper is organized as follows. Section 2 describes the data used in this study and presents the results of its preliminary analysis. Section 3 is devoted to deriving a deterministic model of the IERS C01 series. A refined analysis of the IERS C01 PM series using the SSA technique is performed in Section 4. Section 5 sums up the results obtained in this work. Concluding remarks on this study are given in the final Section.
@@ 2 Data description and preliminary analysis
The IERS C01 PM series consists of several parts as described in Table 1. The epoch step in this series is 0.1 yr in 1846—1890, and 0.05 yr starting with 1890.
@@
In this study, the beginning of the IERS C01 series in the date range 1846 to 1899 was used. It lacks $Y_{p}$ data from 1858.9 to 1860.9 inclusive. In total, 21 epochs are missing. The middle gap epoch is 1859.9. The PM data used in this work are shown in Fig. 1
@@
Although the polar motion variations described by the IERS C01 series for 1846–1899 are assumed to consist mainly of AW and CW components, possibly with bias222ftp://hpiers.obspm.fr/iers/eop/eopc01/EOPC01.GUIDE, spectral analysis has shown that in fact the signal has a more complex structure as shown in Fig. 2. These normalized Lomb–Scargle periodograms were computed using the algorithm given by Zechmeister and Kürster (2009) after removing the bias in $X_{p}$ and $Y_{p}$ series clearly visible in Fig. 1. As expected, the main spectral peaks correspond to the CW and AW oscillations. However, these spectra show that the C01 series cannot be accurately modeled by a simple BCA model, and more refined analysis adequately corresponding to the data is needed. Such an analysis of the IERS C01 series will be presented in Section 4.
@@
Figure 2: Normalized Lomb-Scargle periodogram of the IERS C01 series for the interval 1846–1899.
3 Deterministic model
@@
It is not possible to use a straightforward model consisting of two components, CW and AW, because their amplitudes do not remain constant over time in the date range around the gap (Miller, 2011; Zotov et al., 2022). However, these two independent analyses of the IERS C01 series showed that the CW amplitude grew nearly linearly in 1850–1870. Miller (2011) also showed that the AW amplitude is nearly constant in this period with a tiny increase over time. Therefore, it is possible to construct an approximation model for $X_{p}$ and $Y_{p}$ for the period around 1850–1870, consisting of a bias and two oscillations with CW and AW periods and linear amplitude trend:
$Y_{p}=a_{0}+(1+a_{1}^{c}t)\left(a_{c}^{c}\cos\frac{\displaystyle 2\pi t}{\displaystyle P_{c}}+a_{s}^{c}\sin\frac{\displaystyle 2\pi t}{\displaystyle P_{c}}\right)+(1+a_{1}^{a}t)\left(a_{c}^{a}\cos\frac{\displaystyle 2\pi t}{\displaystyle P_{a}}+a_{s}^{a}\sin\frac{\displaystyle 2\pi t}{\displaystyle P_{a}}\right)\,,$ (1)
@@
$\begin{array}[]{rcl}Y_{p}&=&a_{0}+a_{c0}^{c}\cos\frac{\displaystyle 2\pi t}{\displaystyle P_{c}}+a_{s0}^{c}\sin\frac{\displaystyle 2\pi t}{\displaystyle P_{c}}+a_{c1}^{c}t\cos\frac{\displaystyle 2\pi t}{\displaystyle P_{c}}+a_{s1}^{c}t\sin\frac{\displaystyle 2\pi t}{\displaystyle P_{c}}\\[8.61108pt] &{}\hfil&\hskip 11.99998pt+\;a_{c0}^{a}\cos\frac{\displaystyle 2\pi t}{\displaystyle P_{a}}+a_{s0}^{a}\sin\frac{\displaystyle 2\pi t}{\displaystyle P_{a}}+a_{c1}^{a}t\cos\frac{\displaystyle 2\pi t}{\displaystyle P_{a}}+a_{s1}^{a}t\sin\frac{\displaystyle 2\pi t}{\displaystyle P_{a}}\,,\end{array}$ (2)
where $t$ is epoch in years (in fact, with respect to the reference epoch), $P_{c}$=1.19 yr is the period of the CW, and $P_{a}$=1 yr is the period if the AW.
The nine parameters of this model were adjusted on seven intervals of length of 12 to 24 years in 2-year increments centered on the middle epoch of the gap in $Y_{p}$ series, 1859.9. The results of the computation of the filling $Y_{p}$ values are shown in Fig. 3. The differences between 7 variants are within the error of $Y_{p}$ in the IERS C01 series around the gap, which is 0.09′′.
@@
Figure 3: IERS C01 $Y_{p}$ series with filled missed data between 1858.9 and 1860.9 using the BCA model. Plot titles show date intervals used to fit the $Y_{p}$ data to the model.
@@
Generally speaking, an advanced model can be constructed in a similar way with more complicated behavior of trend, and CW and AW amplitudes (and maybe phases), but such an extension does not look reasonable in our case because there is no data enough to fit a many-parameter model. Therefore, analysis based on the BCA model cannot provide a fully satisfactory solution because it relies on a model that only approximately describes the actual pole motion. However, it may be useful for verification of other models. The more reliable solution to the problem to which this study is aimed will be obtained in the next Section using the data-driven SSA technique.
@@ 4 SSA analysis
The SSA method is one of the most powerful tools for analyzing time series with a complex structure (Vautard and Ghil, 1989; Elsner and Tsonis, 1996; Golyandina et al., 2001; Ghil et al., 2002; Golyandina et al., 2018). It is widely used for analyzing geodetic, geophysical, and astronomical time series, in particular, for filling gaps in time series (Kondrashov and Ghil, 2006; Golyandina and Osipov, 2007; Shen et al., 2015; Okhotnikov and Golyandina, 2019; Yi and Sneeuw, 2021; Ji et al., 2023).

### v2

@@ Abstract
The C01 Earth orientation parameters (EOP) series provided by the International Earth Rotation and Reference Systems Service (IERS) is the longest reliable record of the Earth’s rotation. In particular, the polar motion (PM) series beginning from 1846 provides a basis for investigation of the long-term PM variations. However, the pole coordinate $Y_{p}$ in the IERS C01 PM series has a 2-year gap, which makes this series not completely evenly spaced. This paper presents the results of the first attempt to overcome this problem and discusses some ways to fill this gap. Two novel approaches were considered for this purpose: parametric astronomical model consisting of the bias and the Chandler and annual wobbles with linearly changing amplitudes, and data-driven model based on Singular Spectrum Analysis (SSA). Both methods were tested with various options to ensure robust and reliable results. The results obtained by the two methods generally agree within the $Y_{p}$ errors in the IERS C01 series, but the results obtained by the SSA approach can be considered preferable because it is based on a more complete PM model.
@@ 1 Introduction
Time series of the Earth orientation parameters (EOP) are the main source of information for investigation of the Earth’s rotation, in particular the motion of the Earth’s rotation axis in the terrestrial reference frame, which is also called polar motion (PM). Historical overview of the PM observations and analyses can be found in Dick et al. (2000). The PM is a complex phenomenon, which consists of many variational components including trends and (quasi)periodical oscillations with periods from several hours to several decades. In particular, the study of long-term polar motion variations is crucial for expanding our understanding of Earth’s dynamic processes, particularly in terms of the impacts of climate change, changes in Earth’s internal structure and the redistribution of global mass. To study slow PM variations, long time series of the pole coordinates $X_{p}$ and $Y_{p}$ is needed. The longer the time interval covered by EOP series, the lower frequency variations can be detected and investigated.
The international standard of the EOP series is the combined series computed by the International Earth Rotation and Reference Systems Service (IERS) (Bizouard et al., 2019). The IERS C01 EOP series contains the longest PM series derived from astronomical observations lasting from 1846 to now111ftp://hpiers.obspm.fr/iers/eop/eopc01/eopc01.iau2000.1846-now. It is widely used for analysis of the decadal variations in Earth’s rotation (Gaposchkin, 1972; Vondrák, 1988; Rykhlova and Kurbasova, 1990; Kołaczek and Kosek, 1999; Gambis, 2000; Yatskiv, 2000; Schuh et al., 2001; Höpfner, 2004; Guo et al., 2005; Malkin and Miller, 2010; Vityazev et al., 2010; Zotov, 2010; Miller, 2011; Seitz et al., 2012; Zotov and Bizouard, 2012; Beutler et al., 2020; Lopes et al., 2022; Zotov et al., 2022; Yamaguchi and Furuya, 2024).
@@
In this paper, for the first time, to our knowledge, we examine some approaches to address the problem of filling the gap in the IERS C01 series. The challenge is to fill in the missing $Y_{p}$ values in the series while preserving its structure as much as possible. For this purpose, two methods of analysis were considered. The first is the use of a parametric astronomical model, specially developed for this study and consisted of three components: bias, Chandler wobble (CW), and annual wobble (AW) (hereafter referred to as BCA model). The second data-driven model has been constructed using Singular Spectrum Analysis (SSA) (Golyandina et al., 2001). These two methods cover the range of models, from parametric to nonparametric, and allow the treatment of time series that do not have a single fundamental period and are rather short.
The paper is organized as follows. Section 2 describes the data used in this study and presents the results of its preliminary analysis. Section 3 is devoted to deriving a parametric model of the IERS C01 series. A refined analysis of the IERS C01 PM series using the SSA technique is performed in Section 4. Details of SSA analysis are described in Appendix A. Section 5 sums up the results obtained in this work. Concluding remarks on this study are given in the final Section.
@@ 2 Data description and preliminary analysis
The IERS C01 PM series consists of several parts as described in Table 1. The epoch step in this series is 0.1 yr in 1846—1890, and 0.05 yr starting with 1890. In this study, the beginning of the IERS C01 series in the date range 1846 to 1899 was used. Corresponding PM data are shown in Fig. 1
@@
In this study, we have focused on the 19th century part of the IERS C01 PM series, primarily on the period 1846–1889. Although the pole coordinates related to the 19th century have relatively low accuracy compared to PM data obtained after 1900, they are of great importance for investigation of long-term variations the Earth’s rotation. Suffice it to say that as early as the late 19th century Chandler (1894a) separated CW and AW signals, derived the value of the CW period of 428-430 days, which is close to the latter determinations, and discovered an $\approx$66-year period of the CW amplitude variation, which is only about 20% shorter than recently obtained from 180-year PM series. Including the 19th century PM data into modern analyses of the entire IERS C01 PM series provides good evidence for  80-year CW periodicity in amplitude and phase variations, which cannot be reliably estimated using only observations since 1900. All this certainly confirms how useful the 19th century PM series are for studying long-term behavior of the PM.
While the work of Fedorov et al. (1972) is widely available and well documented, the original work Rykhlova (1970a) is less known and available and it makes sense to briefly describe it here. This work was based on the analysis of latitude determinations that were made in the framework of several programs for determination of absolute declinations of stars at three observatories: Greenwich, Pulkovo, and Washington. The Greenwich latitude variations were taken from Chandler (1894b) (years 1836.0–1851.0) and Thackeray (1893) (years 1851.0–1891.5). The Pulkovo latitude variations were taken from Orlov (1961) (years 1842.0–1849.5, 1863.6–1879.0, and 1882.0–1892.0). The Washington latitude variations were derived in Rykhlova (1970b) (years 1845.1–1858.8 and 1861.0–1888.0). All three stations were given equal weights.
The pole coordinates were computed using the formula
$\Delta\varphi=X_{p}\cos\lambda+Y_{p}\sin\lambda\,,$ (1)
where $\lambda$ is the latitude of observatory. For most epochs of the Rykhlova PM series, data from all three observatories were used. In the period when Pulkovo was not observing, the pole coordinates were computed using data from Greenwich and Washington. In the Rykhlova series and, consequently, the IERS C01 PM series there are no $Y_{p}$ data from 1858.9 to 1860.9 inclusive. A total of 21 epochs are missing. The middle epoch of the gap is 1859.9. The gap is explained by absence of observations at Washington observatory during this period. The problem is that the Greenwich and Pulkovo observatories are located at close longitudes, $0^{\circ}$ and $30^{\circ}$, respectively, which is close to direction of $X_{p}$. Therefore, it is impossible to accurately estimate the both pole coordinates using only observations from these two stations. Consequently, only $X_{p}$ values were computed for these epochs. The Washington observatory is located at longitude $-77^{\circ}$, which is close to the $Y_{p}$ direction, and its inclusion in the processing provides a good longitude coverage to reliably determine both pole coordinates.
@@
Although the polar motion variations described by the IERS C01 series for 1846–1899 are assumed to consist mainly of AW and CW components, possibly with bias222ftp://hpiers.obspm.fr/iers/eop/eopc01/EOPC01.GUIDE, spectral analysis has shown that in fact the signal has a more complex structure as shown in Fig. 2. The generalized Lomb-Scargle periodograms for the C01 $X_{p}$ and $Y_{p}$ series were computed according to Zechmeister and Kürster (2009) after removing the bias clearly visible in Fig. 1.
@@
Figure 2: Generalized Lomb-Scargle periodogram of the IERS C01 PM series for the interval 1846–1899, arbitrary units.
3 Parametric model
@@
It is not possible to use a straightforward model consisting of two components, CW and AW, because their amplitudes do not remain constant over time in the date range around the gap (Kołaczek and Kosek, 1999; Yatskiv, 2000; Guo et al., 2005; Miller, 2011; Lopes et al., 2022; Zotov et al., 2022). However, these two independent analyses of the IERS C01 series showed that the CW amplitude grew nearly linearly in 1850–1870. Miller (2011) also showed that the AW amplitude is nearly constant in this period with a tiny increase over time. Therefore, it is possible to construct an approximation model for $X_{p}$ and $Y_{p}$ for the period around 1850–1870, consisting of a bias and two oscillations with CW and AW periods and linear amplitude trend:
$\begin{array}[]{rcl}Y_{p}&=&a_{0}\\ &+&(1+a_{1}^{c}t)\left(a_{c}^{c}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c}}+a_{s}^{c}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c}}\right)\\[8.61108pt] &+&(1+a_{1}^{a}t)\left(a_{c}^{a}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}+a_{s}^{a}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}\right)\,,\end{array}$ (2)
@@
$\begin{array}[]{rcl}Y_{p}&=&a_{0}\\[4.30554pt] &+&a_{c0}^{c}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c}}+a_{s0}^{c}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c}}\\[8.61108pt] &+&a_{c1}^{c}t\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c}}+a_{s1}^{c}t\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c}}\\[8.61108pt] &+&a_{c0}^{a}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}+a_{s0}^{a}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}\\[8.61108pt] &+&a_{c1}^{a}t\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}+a_{s1}^{a}t\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}\,,\end{array}$ (3)
where $t$ is epoch in years, $t_{0}$=1859.9 (the middle epoch of the gap), $P_{c}$=1.19 yr is the period of the CW, and $P_{a}$=1 yr is the period of the AW.
The nine parameters of this model were fitted using the least squares method (LS) over seven intervals of length of 12 to 24 years with a step of 2 years centered at the middle epoch of the gap in $Y_{p}$ series, 1859.9, in total seven variants. Then the filling $Y_{p}$ values were computed using this model for each variant. The results of these computations are shown in Fig. 3 (upper panel).
@@
Figure 3: IERS C01 $Y_{p}$ series with filled missed data between 1858.9 and 1860.9 using the BCA model (upper panel) and BCA2 (bottom panel). Plot titles show date intervals used to fit the $Y_{p}$ data to the model.
The BCA model represented by Eq. 3 corresponds to the PM model with a single dominant CW frequency, which is generally accepted at present. However, the PM spectrum presented in Fig. 2 reveals two peaks in the CW frequency band with periods of 1.181 yr and 1.231 yr. There may be various reasons for this bifurcation of the PM spectrum, such as an inconsistency in the data on which the series is based, or CW phase variations. In any case, it would be interesting to check how this effect may affect the result of the computation of the modeled $Y_{p}$ in the gap.
For this purpose, the BCA2 model was constructed as follows:
$\begin{array}[]{rcl}Y_{p}&=&a_{0}\\[4.30554pt] &+&a_{c0}^{c1}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c1}}+a_{s0}^{c1}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c1}}\\[8.61108pt] &+&a_{c1}^{c1}t\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c1}}+a_{s1}^{c1}t\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c1}}\\[8.61108pt] &+&a_{c0}^{c2}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c2}}+a_{s0}^{c2}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c2}}\\[8.61108pt] &+&a_{c1}^{c2}t\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c2}}+a_{s1}^{c2}t\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{c2}}\\[8.61108pt] &+&a_{c0}^{a}\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}+a_{s0}^{a}\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}\\[8.61108pt] &+&a_{c1}^{a}t\cos\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}+a_{s1}^{a}t\sin\frac{\displaystyle 2\pi(t-t_{0})}{\displaystyle P_{a}}\,,\end{array}$ (4)
where $t$ is epoch in years, $t_{0}$=1859.9 (the middle epoch of the gap), $P_{c1}$=1.181 yr is the period of the first CW component, $P_{c2}$=1.231 yr is the period of the second CW component, and $P_{a}$=1 yr is the period of the AW. In this case, a longer interval of dates is needed to separate the two CW components. The 13 parameters of this model were fitted using the least squares method (LS) over seven intervals of length of 12 to 24 years with a step of 2 years centered at the middle epoch of the gap in $Y_{p}$ series, 1859.9,. The filling $Y_{p}$ values for the BCA2 model are shown in Fig. 3 (bottom panel). They are close to the results obtained with the BCA model.
To evaluate the real accuracy of the PM series approximation by the parametric models, we computed the differences between the modeled and observed PM values for each of seven variants for the epochs before and after the gap. Then we computed the root mean square error (RMSE) and mean absolute error (MAE) for these differences. The same computations were repeated for the $X_{p}$ series, for which we used the original IERS C01 series from which the epochs 1858.9–1860.9 were cut to make it similar to the $Y_{p}$ series. The results of this error analysis are shown in Table 2. Two main observations can be made from this table. First, shorter intervals provide better approximation, as expected. Secondly, the approximation error for the $X_{p}$ series is better than for the $Y_{p}$ series, which corresponds to a higher precision of the $X_{p}$ values with respect to the $Y_{p}$ values given in the IERS C01 series in the 19th century. Although the approximation error for the BCA2 model is better than that for the BCA model, the latter may be more significant from a scientific point of view because a duality of the CW period is doubtful. For this reason, the average of two first BCA variants and two first BCA2 variants, i.e. those obtained for 12-year and 14-year variants, hereafter referred to as BCA model, will be used for further comparisons.
Table 2: Approximation error for the parametric models BCA and BCA2.
Interval of dates $X_{p}$ $Y_{p}$
MAE RMSE MAE RMSE
BCA model
1853.89–1865.91 0.068 0.081 0.067 0.090
1852.89–1866.91 0.066 0.081 0.071 0.095
1851.89–1867.91 0.079 0.099 0.084 0.110
1850.89–1868.91 0.077 0.096 0.083 0.108
1849.89–1869.91 0.078 0.097 0.084 0.111
1848.89–1870.91 0.079 0.098 0.087 0.112
1847.89–1871.91 0.079 0.098 0.085 0.110
BCA2 model
1853.89–1865.91 0.062 0.076 0.060 0.084
1852.89–1866.91 0.064 0.078 0.060 0.085
1851.89–1867.91 0.073 0.091 0.070 0.095
1850.89–1868.91 0.073 0.092 0.076 0.102
1849.89–1869.91 0.071 0.089 0.077 0.102
1848.89–1870.91 0.071 0.088 0.075 0.099
1847.89–1871.91 0.073 0.092 0.076 0.100
@@
Generally speaking, other advanced models can be constructed in a similar way with more complicated behavior of trend, and CW and AW amplitudes (and maybe phases), but such an extension does not look reasonable in our case because there is no data enough to fit a many-parameter model. Therefore, analysis based on the BCA model cannot provide a fully satisfactory solution because it relies on a model that only approximately describes the actual pole motion. However, it may be useful for verification of other models. The more reliable solution to the problem to which this study is aimed will be obtained in the next Section using the data-driven SSA technique.
@@ 4 SSA analysis
The SSA method is one of the most powerful tools for analyzing time series with a complex structure including the PM series (Vautard and Ghil, 1989; Elsner and Tsonis, 1996; Golyandina et al., 2001; Ghil et al., 2002; Vityazev et al., 2010; Miller, 2011; Golyandina and Zhigljavsky, 2020). It is widely used for analyzing geodetic, geophysical, and astronomical time series, in particular, for filling gaps in time series (Kondrashov and Ghil, 2006; Golyandina and Osipov, 2007; Shen et al., 2015; Okhotnikov and Golyandina, 2019; Yi and Sneeuw, 2021; Ji et al., 2023). SSA is an agile method because it can propose a (possibly local) approximation by a model in the form of a finite sum of products of polynomials, exponentials and harmonics. This model is called linear because it consists of signals governed by linear recurrence relations. However, this is a rather large class of signals, and a signal may only approximately satisfy the model. This means that within a two-year time frame, this property of SSA overcomes its limitations. It should be noted that the set of frequencies does not change during the time interval considered, but rather their amplitudes. By varying the window length, SSA can account for amplitude changes even in the case of some non-linearity.